OpenStudy (anonymous):

evaluate the integral x/(sqrt(1-x^4)dx

OpenStudy (anonymous):

f(x)=1/2*arc sin(x^2)

OpenStudy (anonymous):

let u=x^2 then du/dx= 2x

OpenStudy (anonymous):

Well, here: \[ \frac{d}{dx}\text{asin}(x)=\frac{1}{\sqrt{1-x^2}} \]Dunno what's wrong with me, today.

OpenStudy (anonymous):

thank you guys for helping me

OpenStudy (anonymous):

I really appreciate it!!!!